If a company increases its workforce from 10 to 17 workers and as a result the total output changes from 30 units to 44 units, what is the marginal product for the increased workers?


1
Expert's answer
2021-11-26T12:17:10-0500

Marginal product of labor (MPL) is an increase in total output that occurs when labor increases by one unit, but all other costs remain the same.

"MPL=\frac{\Delta Q}{\Delta L}=\frac{44-30}{17-10}=2"